OpenStudy (anonymous):

Let X represent the outcome of a wheel of fortune that has whole numbers 0 to 50 on different slots. What is the expected outcome? A. 50 B. 25 C. 10 D. 0

OpenStudy (anonymous):

\(\frac{1+2+3+...+50}{51}\)

OpenStudy (anonymous):

since \(1+2+3+...+50=\frac{50\times 51}{2}\) you get \(\frac{50}{2}=25\)

OpenStudy (anonymous):

which i am willing to bet, you could have guessed at

OpenStudy (anonymous):

probability is not my thing i cant get the hang of it.